The symmetric form of the equation of the line of intersection of the planes x+y-z = 1 and 2x-3y+z=2 is

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear deepika

planes x+y-z = 1 and 2x-3y+z=2

 

let the direction ratio of the line are a,b,c

so a+b-c =0   and 2a-3b+c =0

so from these two equation

 a/2  = b/3 =c/5

 

now we have to find out the point of intersection of these two planes.

let x=0

so  y-z =1   and   -3y + z =2

 so y=-3/2

   z = -5/2

 

so equation of line become

 (x-0)/2  = (y +3/2)/3  = (z+5/2)/5

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin